OpenStudy (chloe1221):

i will medal. Whats 2+27+98+7865+8

OpenStudy (jjcoolja12):

2+27+98+7865+8 =29+98+7865+8 =127+7865+8 =7992+8 =8000

OpenStudy (chloe1221):

thank you

OpenStudy (jjcoolja12):

anytime